# A generalization of Barannikov-Kontsevich theorem

### v1

@@ 3.2.2 Completion
The natural morphism is a fine resolution:
@@
Let $G$ be any pseudo-coherent $\mathcal{O}_{X}$-module. (See [5, Appendix A] for pseudo-coherent sheaves.) We have $G\otimes^{L}_{\mathcal{O}_{X}}\Omega^{0,q}_{X}\simeq G\otimes_{\mathcal{O}_{X}}\Omega^{0,q}_{X}$. Let $F_{\bullet}\to G$ be a free resolution of $G$. The natural morphisms $F_{\bullet}\to\mathop{\rm Tot}\nolimits(F_{\bullet}\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X})\to G\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}$ are quasi-isomorphisms. Hence, $G\to G\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}$ is a quasi-isomorphism.
@@
For any open subset $U\subset X$, the morphisms $H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}\to H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ are surjective because $\iota_{\lambda}^{\ast}(\mathcal{G})\otimes\Omega^{0,q}_{X}$ is fine. Let $a_{j}^{q+1}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes\Omega^{0,q+1}\bigr{)}$ $(j=1,2,\ldots)$ be sections such that $\overline{\partial}(a_{j}^{q+1})=0$ and $\pi_{j+1}(a_{j+1}^{q+1})=a_{j}^{q+1}$. Let us construct $b_{j}^{q}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j})\otimes\Omega^{0,q}_{X}\bigr{)}$ $(j=1,2,\ldots)$ such that $\overline{\partial}b_{j}^{q}=a_{j}^{q}$ and $\pi_{j+1}(b_{j+1}^{q})=b_{j}^{q}$ inductively on $j$. Suppose that we have already constructed $b_{j}^{q}$. There exists $c_{j+1}^{q}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ such that $\pi_{j+1}(c_{j+1}^{q})=b_{j}^{q}$. We obtain
@@ 3.3.2 The $E_{1}$-degeneration for another filtration $\mathcal{F}_{u}$
Let $\pi_{1}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to X$ denote the projection. Let $\mathcal{C}_{2}^{\bullet}$ denote the Dolbeault resolution, i.e., the associated complex of the double complex

### v2

@@ 3.2.2 Completion
The following natural morphism is a fine resolution:
@@
Let $G$ be any pseudo-coherent $\mathcal{O}_{X}$-module. (See [6, Appendix A] for pseudo-coherent sheaves.) We have $G\otimes^{L}_{\mathcal{O}_{X}}\Omega^{0,q}_{X}\simeq G\otimes_{\mathcal{O}_{X}}\Omega^{0,q}_{X}$. Let $F_{\bullet}\to G$ be a free resolution of $G$. The natural morphisms $F_{\bullet}\to\mathop{\rm Tot}\nolimits(F_{\bullet}\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X})\to G\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}$ are quasi-isomorphisms. Hence, $G\to G\otimes_{\mathcal{O}_{X}}\Omega^{0,\bullet}_{X}$ is a quasi-isomorphism.
@@
For any open subset $U\subset X$, the morphisms $H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}\to H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ are surjective because $\iota_{\lambda}^{\ast}(\mathcal{G})\otimes\Omega^{0,q}_{X}$ is fine. Let $q\in{\mathbb{Z}}_{\geq 0}$. Let $a_{j}^{q+1}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j}\mathcal{G})\otimes\Omega^{0,q+1}\bigr{)}$ $(j=1,2,\ldots)$ be sections such that $\overline{\partial}(a_{j}^{q+1})=0$ and $\pi_{j+1}(a_{j+1}^{q+1})=a_{j}^{q+1}$. Let us construct $b_{j}^{q}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j})\otimes\Omega^{0,q}_{X}\bigr{)}$ $(j=1,2,\ldots)$ such that $\overline{\partial}b_{j}^{q}=a_{j}^{q}$ and $\pi_{j+1}(b_{j+1}^{q})=b_{j}^{q}$ inductively on $j$. Suppose that we have already constructed $b_{j}^{q}$. There exists $c_{j+1}^{q}\in H^{0}\bigl{(}U,(\mathcal{G}/\lambda^{j+1}\mathcal{G})\otimes\Omega^{0,q}_{X}\bigr{)}$ such that $\pi_{j+1}(c_{j+1}^{q})=b_{j}^{q}$. We obtain
@@ 3.3.2 The $E_{1}$-degeneration for another filtration $\mathcal{F}_{u}$
Let $\pi_{1}:{\mathbb{C}}_{\lambda}\times X\times{\mathbb{C}}_{t}\times{\mathbb{C}}_{u}\to X$ denote the projection. Let $\mathcal{C}_{2}^{\bullet}$ denote the Dolbeault resolution in the $X$-direction, i.e., the associated complex of the double complex